# How precisely can we measure the ages of subgiant and giant stars?

### v1

@@ Abstract
Precise stellar ages are fundamental to Galactic archaeology. However, obtaining reliable age estimates and uncertainties for field stars has been a long-standing challenge. We test the fidelity of ages from recent catalogs of giants and subgiants using wide binaries, whose components formed at the same time and thus should have consistent inferred ages. We find that subgiant ages based on spectroscopic metallicities from Xiang & Rix (2022) are generally consistent within their reported uncertainties, implying that fractional uncertainties of are realistically achievable. In contrast, we find that published photometric subgiant ages and red giant ages underestimate true uncertainties by factors of . These results demonstrate that accurate metallicity and -element abundances are essential for precise isochrone ages and establish wide binaries as a powerful, model-independent benchmark for calibrating stellar age measurements in the era of large spectroscopic surveys.
@@ 3 Results
We find that the catalog of Xiang & Rix (2022) produces the most consistent subgiant ages. For the full sample, we find , which is mostly attributed to a single strong outlier with Age 1 = Gyr and Age 2 = Gyr. Removing this outlier reduces the spread to . The outlier also shows a discrepant [Fe/H] between the two stars, which explains the discrepant age estimates222We check that this source is indeed confidently a wide binary, with a chance alignment probability of (El-Badry et al., 2021) at a physical separation of au and distance pc (Appendix A).. After removing this outlier, the ages of wide binary components are statistically consistent within their reported errors. While this is the smallest wide binary sample of the three (), it is the only one for which .
By contrast, both Nataf et al. (2024) and Wang et al. (2023) appear to underestimate their age uncertainties. For Nataf et al. (2024), the full sample shows substantial excess scatter, and even in the Primary Sample, the dispersion remains well above unity (). This suggests that residual systematics – such as photometric blends, biased extinction values, or metallicity misestimation – inflate the true uncertainties beyond the reported statistical errors. For Wang et al. (2023), the discrepancy is larger still: the full giant sample yields , implying that the uncertainties are underestimated by a factor of 5 on average. Unlike in the Xiang & Rix (2022) sample, it is not possible to substantially reduce the scatter here by removing any one outlier. Restricting to SNR spectra somewhat improves the age precision (), but this still implies that the reported uncertainties are underestimated by a factor of .
In summary, for the Nataf et al. (2024) and Wang et al. (2023) samples indicates that their reported age uncertainties are systematically underestimated. The reduction in scatter from the full Nataf et al. (2024) sample to their Primary Sample, and the improved agreement for high-SNR () sources in Wang et al. (2023), show that stringent quality cuts enhance reliability. Nevertheless, even after these cuts, the true uncertainties remain underestimated by factors of . Among the three age catalogs considered, only Xiang & Rix (2022) achieve age estimates that are both precise and empirically validated by wide binaries. Their success likely stems from the use of spectroscopic [Fe/H] and [/Fe] measurements to anchor the isochrone grids for each star. The median fractional age uncertainty in their sample is .
@@ 4.1 Interpreting wide binary Constraints
These limitations imply that while wide binaries provide a useful test of reported uncertainties, they do not fully capture absolute accuracy. Systematic uncertainties associated with stellar evolution models set another floor for absolute age uncertainties (e.g. Ying et al., 2025), which are crucial for cosmological and chemo-dynamical applications where absolute ages matter, such as in calibrating the age of the Universe or the early enrichment timeline. Our results therefore complement, rather than replace, absolute calibration efforts.
@@ 5 Conclusions
Our analysis shows that the subgiant ages of Xiang & Rix (2022) are consistent with their quoted precisions: after removal of a single strong outlier, the scatter between binary components is fully consistent with the reported errors (). In contrast, both the catalogs of Nataf et al. (2024) and Wang et al. (2023) underestimate age uncertainties, with wide binary tests indicating effective errors larger than reported formal errors. Even among their quality-controlled subsamples, we find , demonstrating that additional systematics remain unaccounted for.
A key difference between the subgiant samples of Xiang & Rix (2022) and Nataf et al. (2024) lies in their treatment of metallicity. The former use spectroscopic measurements of both [Fe/H] and (critically) [/Fe], highlighting the importance of accurate chemical abundances for reliable isochrone ages.
This study underscores the utility of wide binaries as a model-independent benchmark for stellar age measurements. Unlike internal validation tests or calibrations to clusters, wide binaries provide a large set of validation experiments across the Galaxy, and most importantly, across a wide range of ages, metallicities, and -abundances, enabling robust empirical assessments of true uncertainties.

### v2

@@ Abstract
Precise stellar ages are fundamental to Galactic archaeology. However, obtaining reliable age estimates and uncertainties for field stars has been a long-standing challenge. We test the fidelity of ages from recent catalogs of giants and subgiants using wide binaries, whose components formed at the same time and thus should have consistent inferred ages. We find that subgiant ages based on spectroscopic metallicities from Xiang & Rix (2022) are generally consistent within their reported uncertainties, implying that fractional uncertainties of are realistically achievable. In contrast, we find that published photometric subgiant ages underestimate true uncertainties by factors of . Spectroscopic age estimates for red giant and red clump stars also show reliable uncertainties, but are generally less precise (). These results demonstrate that accurate metallicity and -element abundances are essential for precise subgiant ages and establish wide binaries as a powerful, model-independent benchmark for calibrating stellar age measurements in the era of large spectroscopic surveys.
@@ 3 Results
We find that the catalog of Xiang & Rix (2022) produces the most consistent subgiant ages. For the full sample, we find , which is mostly attributed to a single strong outlier with Age 1 = Gyr and Age 2 = Gyr. Removing this outlier reduces the spread to , fully consistent with the reported uncertainties. The outlier also shows a discrepant [Fe/H] between the two stars, likely explaining the divergent ages222We check that this source is indeed confidently a wide binary, with a chance alignment probability of (El-Badry et al., 2021) at a physical separation of au and distance pc (Appendix A).. Despite the modest sample size (), these results demonstrate that subgiant ages from Xiang & Rix (2022) are statistically consistent with their quoted uncertainties ().
For Wang et al. (2023), the reported age errors are also consistent with the observed scatter: the full red-giant and red-clump sample yields , improving slightly to for spectra with SNR. These results indicate that age precisions of are realistically achievable for giants, though they remain less precise than the age uncertainties inferred for subgiants in Xiang & Rix (2022).
By contrast, the Nataf et al. (2024) catalog appears to underestimate age uncertainties. The full sample shows substantial excess scatter, and even in the Primary Sample, the dispersion remains well above unity (). This suggests that residual systematics – such as photometric blends, biased extinction values, or metallicity misestimation – inflate the true uncertainties beyond the reported statistical errors. It also highlights the importance of accurate abundance measurements for subgiant ages.
In summary, for the Xiang & Rix (2022) and Wang et al. (2023) samples indicates that their reported age uncertainties are accurate, while for the Nataf et al. (2024) sample indicates that their errors are underestimated by a factor of . The reduction in scatter from the full Nataf et al. (2024) sample to their Primary Sample, and the improved agreement for high-SNR () sources in Wang et al. (2023), show that stringent quality cuts enhance reliability. Among the three age catalogs considered, Xiang & Rix (2022) achieve the most precise age estimates that also have empirically validated errors, as determined by wide binaries. The median fractional age uncertainty in their sample is . While the Wang et al. (2023) sample also has consistent errors for red giant and red clump stars, their median uncertainties are () for the total (SNR) sample. The success of Xiang & Rix (2022) reflects both the greater intrinsic age sensitivity of subgiants and the use of spectroscopic [Fe/H] and [/Fe] measurements to anchor their isochrone modeling.
@@ 4.1 Interpreting wide binary Constraints
These limitations imply that while wide binaries provide a useful test of reported uncertainties, they do not fully capture absolute accuracy. Systematic uncertainties associated with stellar evolution models set another floor for absolute age uncertainties (e.g. Morales et al., 2025; Ying et al., 2025), which are crucial for cosmological and chemo-dynamical applications where absolute ages matter, such as in calibrating the age of the Universe or the early enrichment timeline. Our results therefore complement, rather than replace, absolute calibration efforts.
@@ 5 Conclusions
Our analysis shows that the subgiant ages of Xiang & Rix (2022) and the red giant/red clump ages from Wang et al. (2023) are consistent with their quoted precisions (). The median fractional age uncertainty in the Xiang & Rix (2022) and Wang et al. (2023) sample is and , respectively.
In contrast, subgiant ages from the Nataf et al. (2024) catalog exhibit underestimated age uncertainties, with wide binary tests indicating effective errors larger than reported formal errors. Even among their quality-controlled subsample, we find , demonstrating that additional systematics remain unaccounted for.
A key difference between the subgiant samples of Xiang & Rix (2022) and Nataf et al. (2024) lies in their treatment of chemical abundances. The former use spectroscopic measurements of both [Fe/H] and (critically) [/Fe], highlighting the importance of accurate chemical abundances for reliable isochrone ages. For Galactic archaeology applications that demand high-precision ages, the Xiang & Rix (2022) subgiant catalog currently offers the most precise and empirically validated ages among those considered here.
Beyond testing reported age precisions, this study also underscores the utility of wide binaries as a model-independent benchmark for stellar age measurements. Unlike internal validation tests or calibrations to clusters, wide binaries provide a large set of validation experiments across the Galaxy, and most importantly, across a wide range of ages, metallicities, and -abundances, enabling robust empirical assessments of true uncertainties.